OpenStudy (anonymous):

Find the sum of 2√6 and √24 A. 4√6 B. 2√30 C. 2√6 4√3

OpenStudy (anonymous):

Sos

OpenStudy (anonymous):

c

OpenStudy (anonymous):

jk change sqrt 24 to sqrt4 X sqrt 6 which would be 2 sqrt 6 so add them and u get A-- but this is cheating -- dont get cuaght

OpenStudy (anonymous):

its A, \[2\sqrt{6} + \sqrt{24} = 2\sqrt{6} + \sqrt{4}\sqrt{6} = 2\sqrt{6}+2\sqrt{6} = 4\sqrt{6}\]

OpenStudy (anonymous):

2sqrt (6) + 2ssqrt(6) givesw 4 sqrt6)

OpenStudy (savvy):

24 can be written as 4X6...... so \[\sqrt{24} = \sqrt{4\times6} = \sqrt{4}\times \sqrt{6}=2\times \sqrt{6}\] so add it....o get 4sqrt.6

OpenStudy (anonymous):

thanx